if i am enlargeing a birthday candle by 5 times its size, and the diameter is 1/4 in and the height is 2 1/2 inches how do i find the volume?

Found 2 solutions by Alan3354, checkley75:
Answer by Alan3354(69443)

if i am enlargeing a birthday candle by 5 times its size, and the diameter is 1/4 in and the height is 2 1/2 inches how do i find the volume?
------------------
What do you mean by size? And, are you making it 5 times as big, or adding 5 times what its present "size" is?

Answer by checkley75(3666)

V=BASE*HEIGHT
V=(PI*D)*2.5=3.14*.25*2.5=1.9625 IN^3 IS THE ORIGINAL VOLUME.
5*1.9625=9.8125 IN^3 IS THE ENLARGED VOLUME.